OpenStudy (anonymous):

The sum of two numbers is 82. Their difference is 24. Write a system of equations that describes this situation. Solve by elimination to find the two numbers.

OpenStudy (cwrw238):

x + y = 82 x - y = 24 now if you add these 2 equations you eliminate y then solve for x and finally plug this value of x into the first equation and solve for y

OpenStudy (anonymous):

thank u!

OpenStudy (cwrw238):

yw

OpenStudy (anonymous):

fyi. i got this wrong, but idk why

OpenStudy (cwrw238):

what answers did you get/

OpenStudy (anonymous):

idk what it is, it wont tell me which answer is correct

OpenStudy (cwrw238):

adding the equations gives 2x = 106 x = 53 plug into ist equation: 53 + y = 82 y = 82-53 = 31